OpenStudy (anonymous):

sq rt (4x^2) * sq rt (28x)

OpenStudy (anonymous):

\[\sqrt {4x^2} * \sqrt {28x}\]look at the 4x^2 term first. What can that be simplified to?

OpenStudy (anonymous):

4x because the square root and the power cancel

OpenStudy (anonymous):

There is a property where \[\large (ab)^x=a^xb^x\]So the 1/2 power has to be applied to each term under the radical

OpenStudy (anonymous):

half?

OpenStudy (anonymous):

\[\huge \sqrt x=x^{1/2}\]

OpenStudy (anonymous):

that is what the radical means

OpenStudy (anonymous):

\[\large \sqrt {4x^2}=(4x^2)^{1/2}=4^{1/2}x^{2*1/2}\]

OpenStudy (anonymous):

\[\sqrt{4x^2}=\sqrt{4}\times \sqrt{x^2}=2x\]

OpenStudy (anonymous):

2x*28x

OpenStudy (anonymous):

28x is still under the radical tho. We haven't simplified that part yet

OpenStudy (anonymous):

Now sqrt(x) can't be simplified but we can do something for sqrt(28) we can split it up into its factors, one being a perfect square so that we can simplify it. \[\sqrt{28}=\sqrt{4*7}=\sqrt4\sqrt7=2\sqrt7\]

OpenStudy (anonymous):

\[\large 2x*2\sqrt{7x}=4x \sqrt{7x}=4x^{3/2}\sqrt7\]

OpenStudy (anonymous):

where did the 3/2 come from?

OpenStudy (anonymous):

\[\huge x^ax^b=x^{a+b}\]